The figures below are several s-t graphs where s is the position of a particle in rectilinear motion and t is time. Rank the graphs from greatest to least amount of particle displacement over the time interval 0 to 3 seconds. Indicate ties as well 20 20 20 20 10 10 10 20 20 10 taken from Conceptual Dynamics by Kirstie Platenberg and Richard Hill
Answer #1

for A: Si = 20 and Sf = 20

displacement = delta(S) = Sf - Si = 0

for B: Si = 10, Sf = 10

displacement = 10 - 10 = 0


For C: Si = 20 , Sf = 0

displacement = 0 - 20 = - 20 m

magnitude = 20 units


For D: Si = 0 , Sf = 20

displacement = 20 - 0 = 20 units


For E: Si = 10 , Sf = 20

displacement = 20 - 10 = 10 units


For F: Si = 0 , Sf = 0

displacement = 0


ans: C = D > E > A = B = F